M 333L 52980 Fall ’97 T Th 12:30-2:00 RLM 7.120

Instructor: C. Friedman RLM 11.140 471-5161 office hrs: T Th 2:30-3:30pm or by appt.

Text and Syllabus: Modern Geometrys The Analytic Approach , Michael Henle, Prentice Hall 1997 Syllabus: Chapters I, II, VII, perhaps part of III.

Prerequisites: M408D

Exams: We will have exams on 9/30 (class 10;Tues.), 11/4 (class 20; Tues.), 12/4 (Class 28; Thurs.). (Final Exam is scheduled for Mon. 12/15, 9-12am, but I am not planning to give one. You should keep this time open, however.)

* Exam grades will be scaled (if the unscaled average is ≤ 75) to some value in the 75 to 80 range. ** Makeups will be given very reluctantly.

Homework: Homework problems will be assigned each Tuesday in class and will be due the next Tuesday. A serious effort should be made on these exercises. It will probably be difficult to completely learn the course material without working problems; also exam questions may bear some resemblance to homework exercises. Additionally, all homework together will be counted as a grade equivalent to one exam grade.

Final Grades: In computing a final grade, I will count the three tests and one total homework grade equally. (So there will be four grades to average.)